OpenStudy (anonymous):

Jeremy is building a model of a square project. The side length of the square base is 12 inches and the slant height of the pyramid is 20 inches. What is the surface area of the model pyramid?

OpenStudy (anonymous):

can someone help me out real quick please

OpenStudy (anonymous):

how many triangular faces are on the pyramid? what does surface area mean? what is the area of a square? what is the area of a triangle?

OpenStudy (anonymous):

please answer these questions

OpenStudy (anonymous):

a. i think four b. isnt that like the bottom part of the pyramid? c.and the area is that the 20 inches?

OpenStudy (goformit100):

Yes it is corect.

OpenStudy (anonymous):

a. a pyramid with a square base has 4 triangular faces b. the surface area of a object is the sum of area of all faces in this case, the surface area of the pyramid would be the area of the square base plus the area of the 4 triangular faces c. no d. blank

OpenStudy (anonymous):

the area of a square can be given by the equation \[A_\square= s^2 = s * s\] the area of a triangle can be given by the equation \[A _\triangle = 1/2* b *h\]

OpenStudy (anonymous):

any questions or comments?

OpenStudy (anonymous):

so for the area of the square it would be.. A=12^2=12*12?

OpenStudy (anonymous):

yes 12*12 =?

OpenStudy (anonymous):

i typed it in 2 ways in case you didnt know that the 2 meant

OpenStudy (anonymous):

oh okay so it would be 12*12=144?

OpenStudy (anonymous):

yes the area of the square base would be 144 inches

OpenStudy (anonymous):

sorry inches squared

OpenStudy (anonymous):

or 12^2=144inches.. oooh okay! and for the last one it would be 1/2*12*20?

OpenStudy (anonymous):

120 inches?

OpenStudy (anonymous):

yes the area of the triangular face would be 120 in^2 now since there are 4 triangular faces, so the total area of the triangular faces would be?

OpenStudy (anonymous):

um... 264..??? i added the 120 and 144

OpenStudy (anonymous):

ok what you have so far is the area of 1 triangular face and the square base you're missing 3 triangular faces

OpenStudy (anonymous):

4 triangular faces + square base 120+120+120+120+144 = ?

OpenStudy (anonymous):

oh 624?

OpenStudy (anonymous):

do you understand what surface area means now?

OpenStudy (anonymous):

um.. i think so :) Thank you i appreciate it!

OpenStudy (anonymous):

also i suggest memorizing the equation for the area of the triangle and the equation for the area of hte square

OpenStudy (anonymous):

Okay i will do that and write it down also :) thanks